OpenStudy (anonymous):

Solve the given equation. Check your solution. 3|2a + 7| =3a + 12

OpenStudy (anonymous):

3 (+2a+7) = 3a+12 , 3(-2a - 7) = 3a+12

OpenStudy (anonymous):

6a + 21 = 3a+12 , -6a -21 = 3a +12 6a -3a = 12-21 , -6a-3a = 12+21 3a = -9 , -9a = 33 a = -9 /3 , a = 33/-9 a = -3 , a = -11/3